OpenStudy (anonymous):

multiply and simplify: 1/2√ 2(√ 6+1/2√ 2) help!

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

1sqr2 x sqr2 = 1(sqr4) = 2 1/2 sqr2(sqr6+2) 1/2 (sqr12+2sqr(2) = .5sqr12+sqr2 =1.732+1.414 =(about) 3.146213

OpenStudy (anonymous):

thanks